# PlainQAFact: Automatic Factuality Evaluation Metric for Biomedical Plain Language Summaries Generation

### v1

@@ 5 Experiments
Since the classifier is trained on the PlainFact training set and its test set is limited to 20 scientific-PLS pairs, we conduct experiments on both PlainFact and the CELLS test set, excluding PlainFact to prevent data leakage.
@@ 5.1 Experiment Settings
Detailed model selections for each module are in §4. Prompts for GPT-4o OpenAI (2024) and Llama 3.1 8B Instruct Meta (2024) are in App. C. For the QG model, we fine-tune the BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs. In MedCPT retrieval, we set k (the number of retrieved snippets) to 3, ensuring the retrieved information remains within a short context window for the subsequent QA process (details in App. D).
@@ 5.2 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work Joseph et al. (2024): (1) Dependency-Arc Entailment (DAE) Goyal and Durrett (2021) is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore Zha et al. (2023) is a model-based factuality metric using RoBERTa Liu (2019). It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC Laban et al. (2022) is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval Fabbri et al. (2022a) is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval Scialom et al. (2021a) is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 6 Results and Analysis
We first analyze the impact of each component incorporated into PlainQAFact and conduct additional ablation studies to assess retrieval effectiveness on PlainFact6.1). Next, we evaluate PlainQAFact on two datasets, comparing it with five widely used factuality metrics (§6.2). Finally, we perform an error analysis to investigate the failure cases of PlainQAFact6.5).
Component Method Choice Benchmark
PlainQAFact 56.0 (0.02)
Classifier Learned classifier [^5] We acknowledge that applying the learned classifier to its training data may introduce data leakage. To further validate PlainQAFact, we evaluate its performance on the CELLS dataset, as shown in Figure 2. -
GPT-4o 56.1 (0.02)
No (retrieve for all) 58.4* (0.03)
Answer Extraction Llama3.1 8B Instruct -
GPT-4o 56.8 (0.05)
PromptRank 54.1*
NP Chunks 53.9*
Retrieval Source Abs + TB + SP -
TB + SP 48.5* (0.02)
Abs (no retrieval) 50.6* (0.03)
Answer Overlap BERTScore -
LERC 28.9* (0.03)
Granular Level Sentence -
Summary 55.8 (0.06)
@@
Table 2: Ablation study on PlainFact, evaluating the impact of individual components in PlainQAFact. The first row represents our proposed combination. Ablations are conducted by replacing one component at a time while keeping all others fixed to match the best overall model. Abs: abstract; TB: Textbooks; SP: StatPearls. Using the first row as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
@@ 6.1 Ablation Study
Table 2 is the sentence-level ablation study of PlainQAFact where each component of our evaluation metric is modified individually. See App. E for the summary-level evaluation. Additional analysis of retrieval resources is presented in Table 3.
Learned Classifier
Evaluating all PLS sentences without classification and retrieving for all improves factuality assessment but significantly increase evaluation time. Retrieval for every sentence extends processing time from 2 hours and 40 minutes with our learned classifier to 6 hours and 30 minutes without it for 2,740 evaluation examples. This highlights the efficiency gains of incorporating a classifier in PLS factuality evaluation. Notably, the GPT-4o classifier achieves results comparable to our learned classifier trained on expert-annotated data, demonstrating the effectiveness of LLMs as alternative classifiers. This extends the applicability of PlainQAFact to factuality evaluation in other domains, as LLMs can serve as classifiers without requiring costly human annotations.
Llama for Gold Answer Extraction
The Llama-based answer extraction method outperforms NP Chunks by 2.1% by providing context-aware, flexible boundaries and robust answer selection Yu et al. (2024), making it more effective for factuality evaluation. Although GPT-4o achieves slightly better performance, the improvement is marginal, making Llama a more cost-effective alternative.
Sentence Type Source Reference Benchmark
Simplification Abs 63.3 (0.05)
Explanation Abs 30.5* (0.03)
TB 17.0* (0.02)
SP 18.4* (0.03)
TB + SP 25.2* (0.02)
Abs + TB 39.8* (0.08)
Abs + SP 40.1* (0.03)
Abs + TB + SP 44.6 (0.06)
Full PLS Abs 50.6* (0.03)
TB 45.4* (0.02)
SP 45.9* (0.03)
TB + SP 48.5* (0.02)
Abs + TB 54.1* (0.02)
Abs + SP 54.3* (0.01)
Abs + TB + SP 56.0 (0.02)
@@
Table 3: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact. The learned classifier categorizes sentences as source simplification (1,685) or elaborativeexplanation (1,055). Abs: abstracts; TB: Textbooks; SP: StatPearls. Using the Abs+TB+SP as the baseline, * indicates a statistically significant difference based on a paired t-test ($p<0.05$). We run each setting for five times and report the standard deviation (std.) in the brackets.
Combined Domain Resources for Retrieval
The best performance is achieved when abstracts, medical Textbooks, and StatPearls are combined as the source document. In contrast, excluding the source abstract for QA-based factuality evaluation achieves the worst performance, which is expected given that abstracts contain most of the key information. We further examine the impact of retrieval resources on different information types. As shown in Table 3, for elaborative explanation sentences, using abstracts alone leads to significantly lower scores compared to source simplification sentences, indicating that abstracts alone are insufficient for verifying external elaborative explanation content. Incorporating StatPearls alongside abstracts substantially improves performance for both elaborative explanation and full PLS, underscoring the importance of integrating high-quality domain-specific resources for factuality verification.
BERTScore for Answer Overlap
@@
Although LERC performs well in QAFactEval, BERTScore achieves significantly better results for PLS factuality evaluation, improving performance by 27.1% on PlainFact. This disparity arises from their fundamental differences: LERC is a learned metric trained on human factual consistency scores, whereas BERTScore relies on embedding-based similarity. Given that PlainQAFact evaluates factuality by comparing generated answers with extracted gold answers from the source, BERTScore’s ability to measure fine-grained semantic similarity makes it more effective choice.
Sentence-level Evaluation
Sentence-level evaluation slightly outperforms summary-level evaluation, primarily due to differences in input context for the QG model. In sentence-level evaluation, the QG model processes one PLS sentence at a time, enabling more targeted and precise factual verification. In contrast, summary-level evaluation takes the entire PLS as input, introducing additional contextual dependencies that can lead to factual inconsistencies. This further confirms the effectiveness of finer granularity in PLS evaluation.
Figure 2: Scaled factuality accuracy on PlainFact and the CELLS test set. PlainQAFact significantly outperforms all other metrics on both datasets, except for AlignScore on PlainFact ($p<0.05$, paired t-test). We compute the std. by running the experiment five times for both our method and Llama 3.1.
6.2 PlainQAFact Achieves the Best Performance in PLS Factuality Evaluation
Figure 2 compares existing factuality metrics with PlainQAFact on the PlainFact and CELLS datasets. To ensure a fair comparison, we scale factuality accuracy to a range of 0 to 1, as QAFactEval scores range from 1 to 5. To further assess generalizability, we evaluate on 200 PLS-abstract pairs randomly sampled from the CELLS test set, sourced from journals outside PlainFact. Results indicate that AlignScore’s effectiveness depends on the proportion of elaborative versus simplification sentences. When fewer external elaborative sentences are present, its performance approaches that of PlainQAFact. However, as elaborative content increases, PlainQAFact consistently outperforms AlignScore, demonstrating its ability to better handle elaborative information absent from the source abstract.
Compared to QA-based metrics such as QAFactEval and QuestEval, PlainQAFact benefits from incorporating retrieval module, which enables a more comprehensive factuality assessment for elaborative explanations. DAE consistently performs the worst in PlainFact settings, highlighting the limitations of dependency-based factuality evaluation for PLS. Additionally, we assess the performance of Llama 3.1 as a judge for factuality evaluation. In summary-level evaluations, Llama 3.1 demonstrates superior performance across both the PlainFact and CELLS datasets, showing its robust capability for context-aware factuality assessment in plain language summarization tasks. However, its performance at the sentence level is inconsistent, while it slightly outperforms PlainQAFact on the CELLS dataset, it decreases by 16.5% on the PlainFact dataset (see App. C for the prompt used with Llama 3.1).
Figure 3: Scaled factuality accuracy on human-annotated elaborative explanation sentences from PlainFact. PlainQAFact significantly outperforms all other metrics on both evaluation levels ($p<0.05$, paired t-test). We compute the std. by running the experiment five times for both our method and Llama 3.1.
@@ 6.2 PlainQAFact Achieves the Best Performance in PLS Factuality Evaluation
Misclassification Plain Language Sentence: This review looked at how well the methods worked to prevent pregnancy[extracted ${}_{\textit{[extracted answer]}}$, if they caused bleeding problems, if women used them as prescribed, and how safe they were. Lopez et al. (2013)
@@
Scientific Abstract: To compare the contraceptive effectiveness, cycle control, compliance (adherence), and safety of the contraceptive patch or the vaginal ring versus combination oral contraceptives (COCs)… Lopez et al. (2013)
@@
QA Answer: cycle control
Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive[extracted ${}_{\textit{[extracted answer]}}$ square that is applied to the skin. Lopez et al. (2013)
Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting. In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…Lopez et al. (2013)
@@
QA Answer: Nonstick dressing Petrolatum-infused gauze strip
Unanswerable Question Plain Language Sentence: Government[extracted ${}_{\textit{[extracted answer]}}$ officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term. von Philipsborn et al. (2019)
Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…von Philipsborn et al. (2019)
@@
QA Answer: long-term effects
Table 4: Error analysis of PlainQAFact with model intermediate outputs. Each plain language sentence is extracted from the PLS, the QG model generates questions based on the extracted answer and PLS sentence, and "Retrieved Knowledge" refers to the context used by the QA model for answer extraction.
6.3 Explanation-Only Evaluation
We evaluate PlainQAFact on human-annotated explanation-only sentences from PlainFact (see Figure 3), conducting both sentence- and summary-level evaluations. For summary-level setting, we aggregate all explanation sentences from each PLS to form a new summary, resulting in 196 summaries in total (4 PLS are labeled as all simplification sentences.). Results show PlainQAFact outperforms all other factuality metrics in assessing explanation information at both evaluation levels, surpassing the second-best metric by 12.2% at the sentence-level and 10.5% at the summary-level. These findings highlight PlainQAFact is effective in evaluating the factuality of elaborative explanations, particularly when external knowledge is required for verification beyond the scientific abstracts. While Llama 3.1 demonstrates superior performance at the summary-level evaluation, PlainQAFact emerges as the best performer in sentence-level setting. Compared to the results in Figure 2, AlignScore performs similarly to PlainQAFact overall on PlainFact but struggles with external sentences, where its performance drops significantly. Thus, PlainQAFact proves to be the most suitable method for the factuality evaluation of PLS.
Figure 4: Evaluation time comparison of PlainQAFact. We compare both sentence- and summary-level evaluation time over four settings: without information retrieval, PlainQAFact, and using GPT-4o as the classifier and answer extractor. We report the percentage reduction in time for summary-level evaluation given the sentence-level running time.
6.4 Evaluation Time
Figure 4 shows the evaluation times for PlainQAFact across both sentence-level and summary-level settings using PlainFact. The percentage scores indicate the reduction in evaluation time when transitioning from sentence-level to summary-level analysis. Our findings reveal that summary-level evaluation consistently reduces evaluation time across all cases. Moreover, employing PlainQAFact without MedRAG-based retrieval results in a 33.3% reduction in evaluation time for sentence-level analysis. Although integrating GPT-4o as both the classifier and answer extractor slightly improves the factuality evaluation performance (shown in Table 2), it also increases the evaluation time, particularly when GPT-4o is used as the classifier.

### v2

@@ 4 Experiments
We conduct main experiments on three publicly available datasets, including PlainFact, CELLS [28], and FareBio [47]. To verify the effectiveness of our proposed PlainQAFact on PLS tasks, we compare with five widely used factual consistency evaluation metrics and two LLM-based evaluators.
4.1 Datasets
As introduced in Sec 3.1, we create PlainFact with fine-grained sentence-level annotations regarding the factuality types. Additionally, to test the generalizability of our pre-trained classifier, we also collect another subset from the test set of CELLS dataset [28], gathering 200 summary-abstract pairs. Similar to PlainFact3.1.1), we also conduct sentence-level perturbation for each plain language summary in our collected CELLS dataset. Therefore, we have 200 factual and 200 non-factual summary-abstract pairs for both PlainFact and CELLS datasets. We also compare PlainQAFact on the FareBio dataset, which contains 175 plain language summaries generated by seven different LLMs, together with the source articles [47]. Specifically, FareBio includes 25 full scientific article examples, each source article is summarized by seven LLMs, resulting 175 plain language summaries in total. We sample plain language sentences that are not both faithful and factual as non-factual sentences, and the rest as factual sentences, resulting in 33 and 174 summaries respectively. We also extract sentences that are labeled as not faithful but factual in their “factual hallucination” label, which represents the plain language sentence that is not from the source article, but still conveying factual information, resulting in 53 valid summaries in total. These sentences are equal to our factual elaborative explanation sentences. Note in FareBio dataset, there are no non-factual simplification sentences. Therefore, we use the same set of non-factual sentences (33 sentences) for main experiment (Table 2) and explanation-only (Figure 2) evaluation.
@@ 4.2 Experiment Settings
The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs. We apply early stopping during training based on validation loss. The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1. The Llama 3.1 model used in AE also uses a temperature of 0.01. For the QG model, we fine-tune BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs. The maximum input length is set to 512 for all models in PlainQAFact. More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in C and  D.
@@ 4.3 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore [19] is a model-based factuality metric using RoBERTa [49]. It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval [50] is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 5 Results and Analysis
We first conduct a pilot study on the FactPICO dataset [9] to investigate the performance of existing factual consistency evaluation metrics on PLS tasks. Our study (G) reveals two major limitations. First, many LLM-generated plain language summaries in FactPICO contain incomplete or incoherent sentences, highlighting the need for a benchmark based on high-quality, human-authored plain language summaries. Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information”, underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts. To address these challenges, we introduce a new expert-annotated dataset, PlainFact, and propose PlainQAFact, a QA-based, retrieval-augmented metric for evaluating factual consistency in plain language summaries.
In this section, we first benchmark PlainQAFact against five widely used automatic factual consistency evaluation metrics on three PLS datasets (§5.1). We then evaluate its performance in an explanation-only setting using PlainFact5.2), followed by an ablation study to assess the contribution of each component (§5.3). Next, we conduct an error analysis to examine failure cases and highlight directions for further improvement (§5.4).
Datasets Metrics Kendall’s $\tau$ Pearson AUC-ROC
CELLS Llama 3.1 57.5 42.1 83.4
GPT-4o 70.0 75.4 99.3
QAFactEval 61.6 70.6 93.5
QuestEval 23.2 28.3 66.4*
SummaC 24.8 29.8 67.5*
AlignScore 56.0 66.2 89.6
DAE 14.4 14.4 55.0*
PlainQAFact 62.1 75.2 93.8
FareBio Llama 3.1 40.5 38.9 83.0
GPT-4o 27.3 35.7 75.7
QAFactEval 32.8 43.9 81.6
QuestEval 37.8 45.2 86.4
SummaC 33.8 32.1 82.5
AlignScore 32.2 47.2 81.0
DAE 8.8 8.8 53.4
PlainQAFact 16.8 35.6 66.1
PlainFact Llama 3.1 60.7 52.6 85.3
GPT-4o 69.8 80.3 99.2
QAFactEval 65.8 79.3 96.4
QuestEval 28.4 34.7 70.1*
SummaC 34.1 42.8 74.1*
AlignScore 60.2 72.6 92.5
DAE 5.3 5.3 51.8*
PlainQAFact 65.7 81.3 96.4
Table 2: Evaluation results of automatic metrics on the CELLS [28], PlainFact, and FareBio [47] datasets. Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC, with eight metric scores compared against human-labeled factuality. The standard deviations (std.) over five runs are: 0.2 (PlainQAFact), 0.5 (Llama 3.1), and 6.0 (GPT-4o) on PlainFact; 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 6.3 (GPT-4o) on CELLS; and 0.1 (PlainQAFact), 2.9 (Llama 3.1), and 17.0 (GPT-4o) on FareBio. * indicates an improvement of our metric over prior work with 95% confidence interval (details in  E). Note the plain language summaries in FareBio are generated by seven LLMs, while CELLS and PlainFact are written by the authors of original articles.
5.1 Main Results
We report rank correlation (Kendall’s $\tau$), linear correlation (Pearson), and discrimination (AUC-ROC [51]) for three datasets (PlainFact, CELLS [28], and FareBio [47]) following previous studies [16, 39, 49] in Table 2. Overall, GPT-4o performs the best on CELLS and PlainFact, but its performance drops on FareBio. DAE is consistently the weakest across all datasets and metrics, suggesting that a dependency-only metric is insufficient for capturing the factual consistency required in PLS tasks.
On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria (Kendall’s $\tau$=62.1, Pearson=75.2, AUC-ROC=93.8). This shows that our approach, first detecting elaborative content and then verifying it with retrieval, works better than existing QA-based metrics, which ask all questions from the source only, and also better than NLI-based evaluation metrics. Similarly, on PlainFact, PlainQAFact achieves the highest Pearson (81.3), surpassing GPT-4o by 1.0 point. It ties AUC-ROC (96.4) score with QAFactEval and has a competitive Kendall’s $\tau$ (65.7). These results on both datasets show that PlainQAFact is reliable for evaluating both factual and non-factual plain language summaries.
On FareBio, no single method performs the best on all criteria. QuestEval achieves the best AUC-ROC (86.4), AlignScore has the best Pearson (47.2), while Llama 3.1 reaches the best Kendall’s $\tau$ (40.5). However, FareBio differs from the other two datasets in several ways: (1) the original labels in FareBio are defined at the sentence level, and a summary is marked as non-factual if any sentence is non-factual; (2) their summaries are generated from full scientific articles rather than abstracts, making it harder for metrics to locate matching source text; (3) the dataset is imbalanced (33 non-factual vs. 174 factual); and (4) All plain language summaries are generated by LLMs and no elaborative explanation annotations are provided for each sentence or summary. These factors likely increase evaluation difficulty on FareBio and impact the performance of multiple metrics, including PlainQAFact and GPT-4o.
In summary, PlainQAFact is a more consistent and effective metric for factual consistency compared to existing automatic approaches. It clearly improves over prior QA-based metrics on CELLS and matches or exceeds them on PlainFact. Although GPT-4o performs better in most of the settings, our PlainQAFact is built entirely on the open-source LLM (i.e., Llama 3.1), ensuring transparency, reproducibility, and accessibility. Its two-step evaluation process, detecting explanations and verifying them with retrieval, helps align metric scores more closely with human judgments in PLS evaluation tasks.
Figure 2: Overall performance on human-annotated elaborative explanation summaries from PlainFact (392 summaries). The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.1, 1.0, and 7.7, respectively based on five runs for each metric. * indicates an improvement of our metric over prior work with 95% confidence interval (details in  E). PlainQAFact significantly outperforms most of the automatic factual consistency evaluation metrics in AUC-ROC. Note that the CELLS dataset does not contain annotations for elaborative explanations. Results of explanation-only evaluation on FactPICO and FareBio are reported in F.
5.2 Explanation-Only Evaluation
To test whether existing evaluation metrics are limited in assessing elaborative explanations in PLS, we evaluate PlainQAFact on human-annotated explanation-only summaries from PlainFact. We select only the sentences labeled as “elaborative explanation” and group them into plain language summaries, resulting in 392 summary–abstract pairs for PlainFact. Note that CELLS does not provide sentence-level annotations for elaborative explanations, so we evaluate it only as a general PLS dataset (§5.1).
As shown in Figure 2, PlainQAFact outperforms all existing automatic metrics across all three criteria. It is 5.7 points higher in Kendall’s $\tau$, 7.5 points higher in Pearson, and 4.0 points higher in AUC-ROC compared with the second-best metric, QAFactEval. These results indicate that PlainQAFact captures the factual consistency of elaborative explanations more effectively than prior automatic metrics. While GPT-4o achieves higher overall scores (Kendall’s $\tau$=54.9, Pearson=60.7, AUC-ROC=88.7), it shows much higher variance (std. 7.7) compared to PlainQAFact (std. 0.1), making it less stable for factual consistency evaluation.
Compared to the results in Table 2, AlignScore is competitive on general PLS tasks but drops on evaluating explanation-only sentences. In contrast, PlainQAFact remains consistently strong when we isolate explanation-only content. PlainQAFact detects added elaborative explanations and checks them with domain knowledge retrieval, which helps when factual consistency depends on facts beyond simple restatement from the source. Since 44% of sentences in our curated benchmark are labeled as elaborative by human annotators, we believe that PlainQAFact is the more suitable and robust factual consistency evaluation metric in this scenario. PlainQAFact also outperforms Llama 3.1 on all three criteria (27.9/28.3/67.0), showing that instruction-tuned LLM judges alone are not enough for reliable evaluation of elaborative content.
Component Method Choice Kendall’s $\tau$ Pearson AUC std.
PlainQAFact 65.7 81.3 96.4 0.2
Classifier Fine-tuned classifier - - - -
GPT-4o 62.8 77.1 94.3* 0.1
No (retrieve for all) 61.2 74.2 93.2* 0.7
Answer Extraction Llama 3.1 - - - -
GPT-4o 66.9 81.9 97.2 0.6
PromptRank 66.3 80.2 96.0 -
Retrieval Source Abs + TB + SP - - - -
TB + SP 63.4 76.8 94.8 0.1
Abs (no retrieval) 62.8 77.1 94.3* 0.1
Answer Overlap BERTScore - - - -
LERC 64.5 75.9 95.5 0.1
Granular Level Sentence - - - -
Summary 49.3 59.6 84.7* 0.5
@@ 5.2 Explanation-Only Evaluation
Table 3: Ablation study of PlainQAFact on PlainFact, analyzing the impact of individual components in PlainQAFact. Abs: abstract; TB: Textbooks; SP: StatPearls. AUC represents AUC-ROC. The first row of each component setting represents our best combination. * indicates an improvement of our metric over prior work with 95% confidence interval. We run each setting for five times and report the standard deviation (std.). Note that only the settings using LLMs may produce fluctuated scores.
@@ 5.3 Ablation Study
PlainQAFact consists of several modules, including a fine-tuned classifier, QA modules, a retrieval function, and an answer overlap evaluation process. To understand the contribution of each component, we conduct an ablation study on PlainFact5.3). Table 3 summarizes the ablation results, where each component of PlainQAFact is individually modified to measure its impact on overall performance.
Fine-tuned Classifier Table 3 shows that removing the classifier and retrieving for every sentence does not improve factual consistency assessment. Performance drops across all criteria (Kendall’s $\tau$ from 65.7 to 61.2; Pearson from 81.3 to 74.2; AUC-ROC from 96.4 to 93.2), and the std. increases to 0.7. This approach also requires notably more computation. Using GPT-4o as a simple classifier provides reasonable results but still underperforms compared to our fine-tuned classifier. These findings support the use of a lightweight fine-tuned classifier to trigger retrieval only when necessary.
Answer Extraction For extracting gold answers from plain language summaries, the LLM-based extractors are better than PromptRank [40]. GPT-4o achieves the best overall scores (Kendall’s $\tau$=66.9, Pearson=81.9, AUC-ROC=97.2), with small gains over our Llama-based method. PromptRank performs better on Kendall’s $\tau$ (increases 0.6), but fails on other two criteria compared to our best combination. Although GPT-4o achieves slightly higher performance, the marginal gain does not justify the increased cost, making Llama 3.1 a more practical and cost-effective choice for large-scale factual consistency evaluation.
Component Method Choice Kendall’s $\tau$ Pearson AUC std.
Source Simplification Abs 63.5 79.0 94.8 0.2
Elaborative Explanation Abs 32.8 38.2 72.8* 0.2
TB 31.2 37.2 70.6* 0.1
SP 33.3 37.8 72.5 1.7
TB + SP 36.7 42.0 75.4 0.1
Abs + TB 40.3 47.0 78.2 0.2
Abs + SP 40.5 47.4 78.3 0.2
Abs + TB + SP 44.2 51.3 81.0 0.1
Full Dataset Abs 62.8 77.1 94.3* 0.1
TB 60.8 73.8 93.0* 0.1
SP 61.3 74.2 93.3* 0.1
TB + SP 63.4 76.8 94.8 0.1
Abs + TB 64.3 79.7 95.4 0.2
Abs + SP 65.0 80.3 95.9 0.2
Abs + TB + SP 65.7 81.3 96.4 0.2
@@
Table 4: Ablation study results on the PlainFact, evaluating how different retrieval sources affect various information types using PlainQAFact. The fine-tuned classifier categorizes input sentences as source simplification or elaborative explanation. Overall, the numbers of factual summaries that only include source simplification and elaborative explanation sentences are 198 and 191 respectively, and 198 and 194 for non-factual summaries. Abs: abstracts; TB: Textbooks; SP: StatPearls. AUC represents AUC-ROC. * indicates an improvement of Abs + TB + SP over other settings with 95% confidence interval. We run each setting for five times and report the standard deviation (std.) in the brackets.
Combined Domain Resources for Retrieval The highest performance is achieved when combining all three sources (Abs + TB + SP). This suggests that comprehensive retrieval, which includes the original source abstracts, improves overall factual consistency assessment. Compared to using only external sources (TB + SP), performance drops by 2.3/4.5/1.6 across the three criteria. Using only abstracts (i.e., no elaborative retrieval) also lowers performance (2.9/4.2/2.1). To further examine the effectiveness of retrieval for source simplification and elaborative explanation, Table 4 breaks down the results as follows:
(1) Source simplification: These are sentences in plain language summaries classified as “source simplification.” We evaluate summaries containing only simplification sentences, using abstracts as the source. Using abstracts as sole source texts achieves strong performance (63.5/79.0/94.8) across all three criteria.
(2) Elaborative explanation: For summaries containing only explanation sentences, abstracts alone are not sufficient (32.8/38.2/72.8), performing much worse than in simplification cases. Adding abstracts with external sources improves performance, with the best results obtained by combining all three sources (Abs + TB + SP).
(3)Full dataset: Retrieval from StatPearls provides better results than retrieval from Textbooks, highlighting the importance of using high-quality, domain-specific knowledge bases for PLS evaluation. Overall, the best combination across all settings is Abs + TB + SP (65.7/81.3/96.4).
These findings show that effective retrieval for factual consistency should include both the source abstract and reliable external medical references, especially when summaries contain elaborative explanations.
@@
Answer Overlap Evaluation BERTScore outperforms the LERC approach introduced in QAFactEval [21], with gains of 1.2/5.4/0.9 across the three evaluation criteria. This performance gap can be attributed to the fundamental differences between the two metrics. LERC is a learned model trained to predict human factual consistency scores, which may reduce its generalizability to new domains or sentence types. In contrast, BERTScore computes embedding-based semantic similarity, allowing it to capture fine-grained semantic overlap between generated and gold answers. Since PlainQAFact evaluates factual consistency by comparing model-generated answers to those extracted from the source, BERTScore’s sensitivity to semantic alignment makes it a more effective and robust choice for this step of the evaluation.
Sentence-level vs. Summary-level Evaluation To further examine the effect of input granularity in factual consistency evaluation, we deactivate the sentence-splitting function and instead pair each plain language summary with an abstract to evaluate factual consistency at the summary level. Our best sentence-level evaluation achieves an AUC-ROC of 96.4, while the summary-level setting performs worse across all three criteria (Kendall’s $\tau$=49.3, Pearson=59.6, AUC-ROC=84.7; std. 0.5). Evaluating at the summary level introduces broader contextual dependencies, which can omit sentence-specific factual errors or create inaccurate entailments. These results suggest that processing one sentence at a time makes QA-based question generation and answering more focused and reduces noise from unrelated context.
Correct Case Plain Language Sentence: Limits to the availability of SSB in schools (e.g. replacing SSBs with water in school cafeterias). [52]
Scientific Abstract: Frequent consumption of excess amounts of sugar-sweetened beverages (SSB) is a risk factor for obesity, type 2 diabetes, cardiovascular disease and dental caries… [52]
Generated Question: Limits to the availability of what in schools?
Retrieved Knowledge: Even though additional data is required to determine the impact of limiting the availability of nutrient-poor or high-sugar goods in schools on obesity, some study results have shown a net-positive result…
Final Answer: nutrient-poor or high-sugar goods
@@
Misclassification Plain Language Sentence: This review looked at how well the methods worked to prevent pregnancy, if they caused bleeding problems, if women used them as prescribed, and how safe they were. [53]
@@
Scientific Abstract: To compare the contraceptive effectiveness, cycle control, compliance (adherence), and safety of the contraceptive patch or the vaginal ring versus combination oral contraceptives (COCs)… [53]
@@
Final Answer: cycle control
Retrieval Quality Plain Language Sentence: The patch is a small, thin, adhesive square that is applied to the skin. [53]
Scientific Abstract: Users of the norelgestromin-containing patch reported more breast discomfort, dysmenorrhea, nausea, and vomiting. In the levonorgestrel-containing patch trial, patch users reported less vomiting, headaches, and fatigue…[53]
@@
Final Answer: Nonstick dressing Petrolatum-infused gauze strip
Unanswerable Question Plain Language Sentence: Government officials, business people and health professionals implementing such measures should work together with researchers to find out more about their effects in the short and long term. [52]
Scientific Abstract: To assess the effects of environmental interventions (excluding taxation) on the consumption of sugar-sweetened beverages and sugar-sweetened milk, diet-related anthropometric measures and health outcomes, and on any reported unintended consequences or adverse outcomes…[52]
@@
Final Answer: long-term effects
Table 5: Error analysis of PlainQAFact with intermediate metric outputs. “Retrieved Knowledge” refers to the source texts retrieved for each plain language sentence. We present correct and failure cases sampled only from elaborative explanation examples (i.e., sentences classified as explanations during evaluation). The correct case has a PlainQAFact score above 0.6, while the failure cases have scores below 0.5.
Color legend: extracted answer, answer origin, correct answer, incorrect answer.